The Mars Pathfinder Mission
Status Reports - Fifth Week

1 August 1997, 6:00 p.m. PDT

A wealth of new information about Martian weather and atmosphere was acquired by the Pathfinder lander on the 27th and 28th days of surface operations in the Ares Vallis outflow channel on Mars.

Data are showing that daily temperatures do not vary much, but minute-by-minute temperatures can fluctuate dramatically, reported Flight Director Rob Manning. Atmospheric pressures can also change significantly within a matter of minutes. Scientists think some of these variations may be caused by small dust devils that can be whipped up by a gust of wind.

The flight team has been taking advantage of longer downlink sessions over the last several Martian days to acquire as much weather data as possible. Regular weather measurements will be limited in the days ahead due to the lander's limited battery power. A two-day shutdown of lander operations to recharge the battery will occur on Sunday and Monday, Aug. 3-4.

Summer temperatures remain in the same range. This morning's low was minus 75 degrees Celsius (minus 103 degrees Fahrenheit), while highs rose to minus 15 degrees Celsius (5 degrees Fahrenheit).
Winds were blowing lightly from the south and swinging around from the northwest during the day.

The rover continued its trek south today after undergoing a temporary delay. Yesterday the vehicle detected a jam in its left front wheel and autonomously backed up to free the pebble from the cleats in its stainless steel wheels. Today Sojourner marched about 4 meters (13 feet) to the south.

"Using waypoints specified the day before, Sojourner stopped, spun about and carefully backed up and onto a small dune named Mermaid today", Manning said. "This afternoon she continued to back up onto the dune and then lowered the alpha proton X-ray spectrometer onto the dune. Tonight the rover will collect elemental spectra of this interesting feature".

The lander and rover returned more than 60 megabits of science and engineering data on Sols 27 and 28. "This data included yet another rover traverse movie and a series of photometric strips that will allow scientists to precisely gauge the optical properties of key features near the lander", Manning added.

On this Martian day, Sol 28, Earthrise occurred at 3:29 a.m. PDT and sunrise followed at 6:32 a.m. PDT. The Earth set at 5:02 p.m. PDT and sunset occurred at 7:20 p.m. PDT.

2 August 1997, 5:00 p.m. PDT

Mars Pathfinder lander and rover operations were curtailed today when the lander downlink session, scheduled to begin at 1:20 p.m. PDT, was not initiated. The cause of the problem is currently unknown, said Mars Pathfinder Mission Manager Richard Cook.

The flight team, led by Flight Director Carl Steiner, subsequently reestablished communications with the lander and received a brief carrier blip at 3:30 p.m. PDT. Additional downlink sessions were not attempted, however, because of the lack of time before the Earth set over the landing site at 5:42 p.m. PDT.

Data on the health of the lander and rover, in addition to other engineering telemetry, will be acquired tomorrow. The flight team expects Sol 30 operations to proceed normally.

Additional information will be posted on the JPL home page at as it becomes available. An audio update is also available by calling 1-800-391.6654.

On this Martian day, Sol 29, Earthrise occurred at 4:09 a.m. PDT and sunrise followed at 7:11 a.m. PDT. The Sun set over the landing site at 7:59 p.m. PDT.

3 August 1997, 7:00 p.m. PDT

Although yesterday's loss of downlink opportunities has not yet been identified, science activities proceeded normally today on the surface of Mars. Today, Sol 30, marks the end of the Mars Pathfinder primary mission, 30 days after the spacecraft landed in an ancient outflow channel called Ares Vallis.

The Imager for Mars Pathfinder (IMP) continued to image the thin Martian atmosphere, the lander's wind socks, the Sun and the rover as it roamed to another destination, said Carl Steiner, Mars Pathfinder flight director.
Acting as a weather station, the Pathfinder lander -- now called the Sagan Memorial Station -- gathered weather data for the 12th consecutive day. Data from yesterday's surface operations had been stored onboard the lander and were downlinked today. Highs on Mars rose to minus 10 degrees Celsius (14 degrees Fahrenheit) today and dipped to minus 70 degrees Celsius (minus 94 degrees Fahrenheit).

The rover finished its soil analysis of Mermaid Dune before heading toward the Rock Garden. An onboard tilt protection circuit caused the rover to shut down after reaching 10 centimeters (0.3 feet) of motion.

"An especially noisy accelerometer had caused this problem in past, but had successfully guarded the rover against excessive tilt", Steiner said. "The rover team thought it prudent to activate this device, even with the possibility of inadvertent shut-down, because of the uneven path to the Rock Garden and the long traverse". Sojourner will resume this traverse on Sol 32, as Pathfinder's extended mission gets under way.

Three downlink sessions were successfully carried out today using the low-gain antenna once and the high-gain antenna for the next two sessions, Steiner said. The operations team, however, was unable to complete its planned downlink of an eighth (octant) of the so-called "super pan" before the end of the day and the beginning of a two-day sleep period for the lander.

"This will be the first time in nearly 240 days that the lander's electronics will be powered off", Steiner said. "At the conclusion of today's activities, all lander electronics, with the exception of a few computer chips that comprise the hybernate circuit, will be powered off to conserve energy through the Martian evening and prolong our waning battery".

The hybernate circuit has been programmed to wake up the lander at 7:30 a.m. local solar time tomorrow. A backup circuit will wake the lander at 8 a.m. if the lander is still asleep. Tomorrow's activities will focus on recharging the battery to the fullest capacity possible. No science experiments are planned.

On this Martian day, Sol 30, Earthrise occurred at 4:49 a.m. PDT and sunrise occurred at 7:51 a.m. PDT. The Earth set over the landing site at 6:22 p.m. PDT and the sunset was at 8:39 p.m. PDT.


For further information, please visit our website at



An audio update on Pathfinder's status can be heard by calling 1-800-391-6654.


Status reports prepared by:
Office of the Flight Operations Manager
Mars Pathfinder Project
NASA Jet Propulsion Laboratory
California Institute of Technology
Pasadena, CA 91109


To subscribe to the mpf-status list, send e-mail to

with the following statement in the body of the message:

subscribe mpf-status


EDT = PDT + 3 hrs. = UTC - 4 hrs. Venezuelan Time = UTC - 4:30 hrs.

Updated: August 4 '97

Best seen with MS Internet Explorer.

Back to the 25 to 31 July Status Reports in ARVAL

Back to the Pathfinder Mission Status Reports in ARVAL